OpenStudy (anonymous):

11. Which is the solution of the equation below, rounded to the nearest tenth? 2.1x – 4.09 = 6.98x + 1.73 (Points : 4) –1.2 –1.1 –0.5 –0.4

OpenStudy (anonymous):

z=9-17 z=-8

OpenStudy (anonymous):

we have to find the value of z so all no.s and constants should be moved to the other side.and on goin to other side...sign chnges.

OpenStudy (anonymous):

z+17=9 -17 -17 z=-8